# arXiv:2602.22038v2[math.AP] 27 Feb 2026

### v2

@@ 3 Appendix
This appendix contains the proof of Theorem 1 together with several auxiliary inequalities used in the previous sections. We begin by establishing Theorem 1, and subsequently derive the additional estimates required throughout the paper.
@@ 3.1 Proof of Theorem 1
3.1.1 Solving (1) for $\sigma=0$
When $\sigma=0$, a solution for (1) in the sense of Theorem 1, with exception of (17), was constructed in Sections 3 and 4 of [48] and [49]. With this in view, we only need to verify that, it is possible to take $\tilde{C}_{3}$ sufficiently small, in order to get that Gaussian decay.
Proof of (17).
By Theorem 3 in [13] along with (12), we obtain there exists $C>0$ such that for all $(t,x)\in[0,T_{1}]\times\mathbb{R}^{d}$
$\bar{\rho}(t,x)\leq C\widetilde{C}_{3}\int_{\mathbb{R}^{d}}\frac{1}{t^{d/2}}\exp{\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)}dy.$ (71)
Now since
$|x-y|^{2}=|x|^{2}-2x\cdot y+|y|^{2}.$
we find
$\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}$ $=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+\widetilde{C}_{3}|y|^{2}+8t|y|^{2}$
$=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+(\widetilde{C}_{3}+8t)|y|^{2}.$ (72)
Additionally, we get
$(\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y.=(\widetilde{C}_{3}+8t)\left(|y|^{2}-2\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}x\cdot y\right).$
and since
$|y|^{2}-2ax\cdot y=|y-ax|^{2}-a^{2}|x|^{2},$
with
$a=\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}.$
we arrive at
$\Big((\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y\Big)=(\widetilde{C}_{3}+8t)|y-ax|^{2}-\frac{\widetilde{C}_{3}^{2}}{\widetilde{C}_{3}+8t}|x|^{2}.$ (73)
Therefore by (72) and (73) we derive
$\exp\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)=\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}\exp\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|y-ax|^{2}\right).$ (74)
and letting z=yaxz=y-ax we have
$\bar{\rho}(x,t)$ $\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\int_{\mathbb{R}^{d}}\exp{\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|z|^{2}\right)}dz$
$\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\left(\frac{\pi 8t\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}\right)^{d/2}$
$\leq C\widetilde{C}_{3}(\pi 8\widetilde{C}_{3})^{d/2}\frac{1}{(\widetilde{C}_{3}+8t)^{d/2}}\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}$ (75)
since
$\int_{\mathbb{R}^{d}}e^{-a|z|^{2}}dz=\left(\frac{\pi}{a}\right)^{d/2}.$
It follows that taking
$\widetilde{C}_{3}<C_{3}\doteq\min\left(\frac{1}{2\pi};\frac{1}{(8\pi)^{d/2}C}\right),$ (76)
for $t\in(0,T]$ such that
$t<T_{1}\doteq\min\left(\frac{\widetilde{C}_{3}}{8};\frac{8(1-2\pi\widetilde{C}_{3})}{2\pi}\right)$ (77)
we end up with
$\bar{\rho}(x,t)\leq 2\exp{(-2\pi|x|^{2})}.$ (78)
3.1.2 Solving (1) for $\sigma\neq 0$
We assume $\tilde{\rho}$ solves the deterministic PDE
$\partial_{t}\tilde{\rho}-\Delta\tilde{\rho}+\nabla\cdot(\tilde{\rho}_{t}K\ast\tilde{\rho}_{t})=0$ (79)
Now we define $\rho(t,x)\doteq\tilde{\rho}(t,x-X_{t})$, where
$dX_{t}=\sigma_{t}dB_{t},$ (80)
Hence, by applying the Itô’s formula, since that $\sigma$ does not depends on spatial variable, we find that $\rho$ solves (1), in the sense of Theorem 1.
@@ 3.2 Inequalities
In this section, we present several inequalities used throughout the text. The first one is the classical Csiszár-Kullback-Pinsker inequality, which is a fundamental result in information theory. It quantifies how control of the relative entropy implies control of the total variation distance.

### v3

@@ 4 Appendix
This appendix contains the proof of Theorem 1 together with several auxiliary inequalities used in the previous sections. We begin by establishing Theorem 1, regarding the existence of a suitable solution of the limiting equation.
@@ 4.1 Proof of Theorem 3
When $\sigma=0$, the equation (1) reads
$\partial_{t}\tilde{\rho}-\Delta\tilde{\rho}+\nabla\cdot(\tilde{\rho}_{t}K\ast\tilde{\rho}_{t})=0,$ (106)
and a solution to (106), in the sense of Theorem 3, was constructed in sections 3 and 4 of [49] and [50]. Now, taking into account (19), we define the following process: $\rho(t,x)\doteq\tilde{\rho}(t,x-X_{t})$.
Hence, by applying the Itô’s formula, since that $\sigma$ does not depends on spatial variable, we find that $\rho$ solves (1), in the sense of Theorem 3, for $\sigma\neq 0$.
@@ 4.2 Inequalities
In this section, we present several inequalities used throughout the text. The first one is the classical Csiszár-Kullback-Pinsker inequality, (see [21]) which is a fundamental result in information theory. It quantifies how control of the relative entropy implies control of the total variation distance.
@@
The next result is a nonlinear Gronwall’s inequality. A proof is provided in Theorem 21 of [17].
Lemma 6.
Let $u$ be a nonnegative function satisfying
$u(t)\leq c(t)+\int_{0}^{t}\big(a(s)\,u(s)+b(s)\,u^{\alpha}(s)\big)\,ds,\qquad c(t)>0,\ \alpha\geq 0,$
where $a(t)$, $b(t)$ and $c(t)$ are integrable and nonnegative functions on $[0,T]$.
For $\alpha=1$,
$u(t)\leq\left(\sup_{t\in[0,T]}c(t)\right)\exp\!\left(\int_{0}^{t}\big(a(s)+b(s)\big)\,ds\right).$ (109)
For $\alpha>1$ and $c$ a constant function, with the following additional hypothesis, for some $h>0$,
$c<\left\{\exp\!\left[(1-\alpha)\int_{0}^{h}a(s)\,ds\right]\right\}^{\frac{1}{\alpha-1}}\left\{(\alpha-1)\int_{0}^{h}b(s)\,ds\right\}^{\frac{1}{1-\alpha}},$ (110)
we get, for $t\in[0,h]$,
$u(t)\leq c\Biggl\{$ $\exp\left[(1-\alpha)\int_{0}^{t}a(s)\,ds\right]$ (111)
$-c^{-1}(\alpha-1)\int_{0}^{t}b(s)\exp\left[(1-\alpha)\int_{s}^{t}a(r)\,dr\right]ds\Biggr\}^{\frac{1}{\alpha-1}}.$